OpenStudy (anonymous):

A huge ocean wave travels a distance of 3.0 x 10^2m in 10 s. calculate the average speed of the wave

OpenStudy (shayaan_mustafa):

\[\large\text{speed}=\frac{\text{distance}}{\text{time}}\] \[\large \text{v}=\frac{\text{s}}{\text{t}}\] \[\large \text{v}=\frac{3.0 \times10^2m}{10s}\] \[\large \text{v}=30\text{m}/\text{s}\]